what is the value for 2-4-6-2+15
Answer by Earlsdon(6103)

Evaluate:
2-4-6-2+15 Work from left-to-right. 2-4 = -2
-2-6-2+15 -2-6 = -8
-8-2+15 -8-2 = -10
-10+15 = 5